Using completing the square what is x? x^2-14x+49 = 64
Answer by usyim88hk(158)

x^2 - 14x + 49 = 64
So we should take out 64 from both sides
x^2 - 14x + 49 - 64 = 64 - 64
x^2 - 14x -15 = 0
Then add 15 to both sides
x^2 - 14x - 15 + 15 = 0 + 15
x^2 - 14x = 15
Now complete the square
x^2 - 14x + (-14/2)^2 = 15 + (-14/2)^2
x^2 - 14x + 49 = 64
Now factor x^2 - 14x + 49
(x-7)^2 = 64
(x-7) = 8 or (x-7) = -8
x-7 = 8 or x-7 = -8
x = 15 or x = -1